Find an equation of a line when it passes throught he points (4/3, -1/2) and
(-2/3, 4).

Answer by checkley71(8403)

(4/3,-1/2)(-2/3,4) gives a slope of (4+1/2)/(-2/3-4/3) or 4.5/-6/3 or 4.5/-2 or
-2.25 for the slope and a y intercept of 2.5 calculated by(y-y1)=m(x-x1) or
y-4=-2.25(X-[-2/3]) OR Y-4=-2.25X-4.5/3
OR Y=-2.25X-1.5+4 OR Y=-2.25X+2.5